# Real-Time Learning of Predictive Dynamic Obstacle Models for Robotic Motion Planning

### v1

@@ III-D Singular Value Hard Thresholding (SVHT)
Given the SVD $Y_{a}=U\Sigma V^{\top}$ and a threshold $\tau$, we approximate the true matrix $X_{a}$ by $\widehat{X}_{a}\;=\;U\Sigma_{\geq\tau}V^{\top}$ where $\Sigma_{\geq\tau}$ is created by setting the elements of $\Sigma$ that are less than $\tau$ to zero.
Gavish and Donoho showed that the Asymptotic Mean Squared Error (AMSE)-optimal threshold for the data singular values depends only on the aspect ratio through the constant $\lambda^{\star}(\beta)$:
@@
with $p_{\beta}(\delta)=0$ outside $\delta_{-},\delta_{+}]$. The median of the MP distribution denoted by $\mu_{\beta}$ depends only on $\beta$ and provides a normalization for estimating the noise level $\sigma$. Taking $\varsigma_{\mathrm{med}}$ as the median of the observed singular values, the the data-driven AMSE–optimal threshold can thus be estimated as:
@@
which remains asymptotically optimal under general white noise (i.i.d., zero mean, unit variance, finite fourth moment) with the same risk as in the Gaussian case [36, Sec. VI].
@@ IV Adaptive Hankel-DMD for Noisy Data
Here, we address the problem outlined in Sec. II. To enable real-time learning and prediction, we consider the use of Hankel-DMD (cf. Sec. III-B). However, applying Hankel-DMD in robotics presents two key challenges:
  • Sensor noise: the data collected from sensors are typically noisy, which contaminates both matrices $H_{0:n-1}^{L}$ and $H_{1:n}^{L}$. This is problematic because, in least-squares schemes such as (4), noise in both data matrices leads to biased or inconsistent estimators [38];
  • Ergodicity and long trajectory data requirements: as noted in Sec. III-B, leveraging the connections between Hankel-DMD and the Koopman operator to capture global dynamics often requires ergodicity and access to long trajectories. These assumptions, however, rarely hold in robotics applications.
In what follows, we first tackle the issue of noise by introducing a denoising scheme for Hankel matrices. We then propose an adaptive sliding-window variation of Hankel-DMD that updates the model in real time as new data arrive. This approach ensures that the model remains accurate within the operating region as the system’s state evolves over time.
@@ IV-A Denoising Hankel Matrices
Consider measurements $\mathcal{B}_{t}=\{x_{t-N+1},\ldots,x_{t}\}$ for buffer size $N$, collected up to time $t$. From this window, we construct a Hankel matrix $H^{\,L}_{t-N+1:t}$ according to Eq. (2). Since the measurements are noisy, our goal is to denoise the Hankel matrix $H^{\,L}_{t-N+1:t}$ via the Cadzow algorithm as described in Sec. III-C. A key step in the Cadzow algorithm is the low-rank projection $\Pi_{r}$ in Eq. (7), which requires knowledge of the effective rank $r$ of the noise-free Hankel matrix. In practice, however, this rank is typically unknown and must be estimated. An arbitrary choice of the rank can cause either over-smoothing (if underestimated) or noise amplification (if overestimated). To avoid this, we adopt a principled, data-driven rank selection strategy based on Singular Value Hard Thresholding (SVHT) (Sec. III-D). However, the SVHT framework assumes i.i.d. noise, an assumption that is generally violated in Hankel matrices due to repeated entries. To address this, we employ Page matrices (Eq. (3)), which partition the measurement buffer into non-overlapping blocks, thereby avoiding redundancy.
@@ IV-A1 Step I: Page-Hankel Rank Transfer
The proof is deferred to the Appendix.
Lemma 1 provides a convenient way to connect the rank of Page and Hankel embedding of noise-free measurement buffers. Before proceeding to rank estimation, we briefly explain that the conditions of Lemma 1 are mild. First, it is worth noting that with a sufficiently high sampling frequency, the system states associated with a given measurement buffer cluster within a small region of the state space, where a local linear model can accurately approximate the system’s behavior. In addition, the condition $\operatorname{span}\{z_{0},Bz_{0},\ldots,(B)^{d-1}z_{0}\}=\mathbb{R}^{n_{z}}$ is generic [^6] If the elements of $A$ (with $B=(A)^{L}$) and $z_{0}$ in Lemma 1 are drawn from a reasonable distribution, the condition holds with probability one. and holds for almost all matrices $B$ and vectors $z_{0}$.
@@ IV-A2 Rank Estimation using SVHT
Now, construct the Page matrix $P^{\,L}_{t-N+1:t}\in\real^{Ln_{x}\times m}$ from the same data sequence used in $H^{\,L}_{t-N+1:t}$. By applying the SVHT formula in Eq.(11), on the Page matrix, we can estimate the singular value threshold $\tau^{*}$ and approximate the effective rank of $P^{\,L}_{t-N+1:t}$ as
@@ V Simulations and Experiments
We apply Algorithm 1 to validate the proposed framework in both simulations and real-world experiments. Our evaluation focuses on (i) signal–to-noise (SNR) separability, (ii) noise variance estimation, (iii) multi-step prediction accuracy for downstream planning and control applications.
@@ V-A Simulation: Noisy Unicycle
Denoising results (Heavy tailed, correlated noise): To probe robustness under non-Gaussian disturbances, we injected temporally correlated AR(1)-Laplace noise (Fig. 2). Despite heavy tails and temporal correlation, the framework remained comparably effective where we observed an SNR gain of $6.9$ dB and an average noise reduction of $54.4\%$, with no visible phase lag relative to the ground truth. While the performance degraded relative to the Gaussian baseline, as expected under correlated, heavy-tailed disturbances, the degradation was gradual. Crucially, the low-rank Hankel structure proved distribution-agnostic i.e., the SVHT adaptively trimmed outlier-inflated singular values without requiring fixed gaussian assumptions on the noise statistic.
@@ V-B System Hardware and Experiments
This section validates the full pipeline on hardware. Our experiment in Fig. 3 is motivated by a naval application. Currently, payloads are loaded into their silos by a ship-board crane while the ship docks in a port, where no ocean swells affect the safety of the loading process. It would be desirable to load these payload units into their silos while the ship is underway, potentially in heavy seas. However, in active sea states, the ship sways about with the ocean swells. Anticipating deck motion would enable the controller to compensate during payload delivery.
The testbed is a scaled crane mounted on an E2M eMove eM6-300-1500 five-DoF Stewart platform that can reproduce wave-induced ship motions. A VN-100 IMU measures linear accelerations and angular velocities of the simulated heaving deck at 30 Hz. Using these measurements, the framework generates 2.0 s i.e., $31$-step horizon forecast.
@@
Given a fixed error-tolerance $\varepsilon=0.04$, we define a violation-duration metric on the prediction error $e_{t}$ as
@@
Together, these results demonstrate that our framework not only denoises real sensor streams but also yields stable and bounded predictions that are suitable for downstream planning and control applications. A real-time demonstration of the moving-base prediction and MPC integration is provided in the supplementary video[42].
@@ VI Conclusion And Future Work
We introduced a Page–Hankel SVHT framework for real-time denoising and short-horizon prediction, yielding stable sliding window Hankel-DMD models from noisy data streams. Simulations (Gaussian and heavy-tailed noise) show variance-stable, distribution-agnostic performance and hardware tests with IMU data produce reliable forecasts that may support downstream control. For future work, we plan to integrate rank-1 updates with exponential smoothing for efficient online adaptation and including tensorized delayed embedding matrices for multi-axis consistency. We also plan to validate our method by integrating the predictions within an MPC framework.

### v2

@@ III-D Singular Value Hard Thresholding (SVHT)
Given the SVD $Y_{a}=U\Sigma V^{\top}$ and a threshold $\tau$, we approximate the true matrix $X_{a}$ by $\widehat{X}_{a}\;=\;U\Sigma_{\geq\tau}V^{\top}$ where $\Sigma_{\geq\tau}$ is formed by zeroing the subthreshold elements of $\Sigma$.
Gavish and Donoho showed that the Asymptotic Mean Squared Error (AMSE)-optimal threshold for the data singular values depends only on the matrix aspect ratio $\beta$ via a constant $\lambda^{\star}(\beta)$:
@@
with $p_{\beta}(\delta)=0$ outside $\delta_{-},\delta_{+}]$. The median of the MP distribution, denoted by $\mu_{\beta}$, depends only on $\beta$ and provides a normalization for estimating the noise level $\sigma$. Taking $\varsigma_{\mathrm{med}}$ as the median of the observed singular values, the data-driven AMSE–optimal threshold can thus be estimated as:
@@
which remains asymptotically optimal under general white noise (i.i.d., zero mean, unit variance, finite fourth moment) with the same risk as in the Gaussian case [14, Sec. VI].
@@ IV Adaptive Hankel-DMD for Noisy Data
We now address the problem outlined in Sec. II. To enable real-time obstacle dynamics learning and motion prediction, we consider the use of Hankel-DMD (cf. Sec. III-B). However, applying Hankel-DMD in real-time robotics applications presents two key challenges:
  • Sensor noise: sensor signals are typically noisy, which contaminates matrices $H_{0:n-1}^{L}$ and $H_{1:n}^{L}$. In least-squares schemes such as (4), noise in both data matrices leads to biased or inconsistent estimators [15].
  • Ergodicity and long trajectory data requirements: as noted in Sec. III-B, one must assume ergodicity and long data sequences to connect Hankel-DMD and the Koopman operator. Such assumptions rarely hold in robotics applications.
To tackle the noise issue, we introduce a Hankel matrix denoising scheme. We then propose an adaptive sliding-window variation of Hankel-DMD that continually updates the model. This approach ensures that the predictive model remains accurate as the object’s state evolves over time.
@@ IV-A Denoising Hankel Matrices
Consider obstacle output measurements $\mathcal{B}_{t}=\{x_{t-N+1},\ldots,x_{t}\}$ for buffer size $N$, collected up to time $t$. We construct a Hankel matrix $H^{\,L}_{t-N+1:t}$ from $\mathcal{B}$ according to Eq. (2). Since the measurements are noisy, we seek to denoise the Hankel matrix $H^{\,L}_{t-N+1:t}$ via Cadzow’s algorithm, as described in Sec. III-C. A key step in this algorithm is the low-rank projection $\Pi_{r}$ in Eq. (7), which requires knowledge of the effective rank $r$ of the noise-free Hankel matrix. In practice, this typically unknown rank must be estimated. An rank estimat error can cause over-smoothing (if underestimated) or noise amplification (if overestimated). To avoid these problems, we adopt a principled, data-driven rank selection strategy based on Singular Value Hard Thresholding (SVHT) (Sec. III-D). However, the SVHT framework assumes i.i.d. noise, an assumption that is violated in Hankel matrices due to repeated entries. To address this, we employ Page matrices (Eq. (3)), whose partition of the measurement buffer into non-overlapping blocks avoids correlation.
@@ IV-A1 Step I: Page-Hankel Rank Transfer
Proof of Lemma 1.
Given the definition of system and the initial condition, one can write $x_{i}=C(A)^{i}z_{0}$ for $i\in\{0,\ldots,N-1\}$. Therefore, one can easily decompose the Page and Hankel matrices as
$P^{L}=EF,\qquad H^{L}=EG,$ (12)
where
$E$ $=[C^{\top},A^{\top}C^{\top},\ldots,\big((A)^{L-1}\big)^{\top}C^{\top}]^{\top}\in\mathbb{R}^{(n_{x}L)\times n_{z}},$
$F$ $=[z_{0},\,Bz_{0},\,(B)^{2}z_{0},\,\ldots,\,(B)^{d-1}z_{0}]\in\mathbb{R}^{n_{z}\times d},$
$G$ $=[z_{0},\,Az_{0},\,(A)^{2}z_{0},\,\ldots,\,(A)^{(d-1)L}z_{0}]\in\mathbb{R}^{n_{z}\times((d-1)L+1)}.$
Note that by the hypothesis, $\operatorname{rank}(F)=n_{z}$, which implies that $F$ has full row rank. Moreover, noting that $B=(A)^{L}$, one can easily see that all columns of $F$ are in the set of columns of $G$; hence, $\operatorname{rank}(G)\geq\operatorname{rank}(F)=n_{z}$. However, note that $G\in\mathbb{R}^{n_{z}\times((d-1)L+1)}$ has more columns than rows; therefore, its rank can be at most $n_{z}$. Hence, one can conclude that both $F$ and $G$ have full row rank and
$\operatorname{rank}(F)=\operatorname{rank}(G)=n_{z}.$ (13)
Moreover, using (12), the rank equality (13), and noting that $\operatorname{rank}(E)\leq n_{z}$ (due to its size) one can write
$\operatorname{rank}(P^{L})\leq\min(\operatorname{rank}(E),\operatorname{rank}(F))=\operatorname{rank}(E),$
$\operatorname{rank}(H^{L})\leq\min(\operatorname{rank}(E),\operatorname{rank}(G))=\operatorname{rank}(E).$ (14)
On the other hand, Sylvester’s rank inequality gives
$\operatorname{rank}(P^{L})=\operatorname{rank}(EF)\geq\operatorname{rank}(E)+\operatorname{rank}(F)-n_{z},$
$\operatorname{rank}(H^{L})=\operatorname{rank}(EG)\geq\operatorname{rank}(E)+\operatorname{rank}(G)-n_{z}.$ (15)
Inequalities (IV-A1)-(IV-A1) in conjunction with (13) lead to $\operatorname{rank}(P^{L})=\operatorname{rank}(H^{L})=\operatorname{rank}(E)$ concluding the proof. ∎
Lemma 1 provides a convenient way to connect the rank of Page and Hankel embedding of noise-free measurement buffers. Before proceeding to rank estimation, we briefly explain that the conditions of Lemma 1 are mild. First, it is worth noting that with a sufficiently high sampling frequency, the system states associated with a given measurement buffer cluster within a small region of the state space, where a local linear model can accurately approximate the system’s behavior [^6] This suggests that the buffer window size and sampling frequency should be chosen jointly to balance fidelity and sensitivity. An excessively large buffer reduces sensitivity to low-rank structures that capture local behavior, whereas an overly small buffer risks discarding expressive information about the system dynamics.. In addition, the condition $\operatorname{span}\{z_{0},Bz_{0},\ldots,(B)^{d-1}z_{0}\}=\mathbb{R}^{n_{z}}$ is generic [^7] If the elements of $A$ (with $B=(A)^{L}$) and $z_{0}$ in Lemma 1 are drawn from a reasonable distribution, the condition holds with probability one. and holds for almost all matrices $B$ and vectors $z_{0}$.
@@ IV-A2 Rank Estimation using SVHT
To construct the Page matrix $P^{\,L}_{t-N+1:t}\in\real^{Ln_{x}\times m}$ from the same data sequence used in $H^{\,L}_{t-N+1:t}$. By applying the SVHT formula in Eq.(11), on the Page matrix, we can estimate the singular value threshold $\tau^{*}$ and approximate the effective rank of $P^{\,L}_{t-N+1:t}$ as
@@ IV-B Online Identification and Multi-Step Prediction
IV-C Comparison with Alternative Online Prediction Methods
Classical online noise filtering approaches, such as the KF and its nonlinear variants[36], assume parametric dynamics and structured noise models. While online Gaussian processes[7, 28] and neural sequence models[25, 10] provide strong predictive capability, they can require significant offline training and careful tuning for online deployment. In contrast, DMD[53] and EDMD[50] are operator-theoretic system identification methods that approximate the Koopman operator from data. These formulations do not account for measurement noise and can exhibit bias under finite, noisy datasets[23, 9], particularly when nonlinear lifting is used[22]. Although noise-aware formulations and regularized variants exist[41], these methods mainly improve dynamical consistency rather than provide a framework for denoising streaming measurements in finite-data regimes. our proposed method explicitly performs structured low-rank denoising prior to short-horizon prediction. Designed for finite, streaming data with limited computational budgets, the framework combines adaptive rank selection with structured matrix embeddings to enable stable real-time performance.
@@ V Simulations and Experiments
We apply Algorithm 1 to validate out framework in simulations and real-world experiments. Our evaluation examines (i) signal–to-noise (SNR) separability, (ii) noise variance estimation, and (iii) multi-step prediction accuracy. Experiments consider open-loop, short-horizon prediction to assess the accuracy, stability and noise robustness of the learned dynamics model. These properties are critical for downstream planning and control frameworks [8].
@@ V-A Simulation: Noisy Unicycle
Denoising results (Heavy tailed, correlated noise): To probe robustness under non-Gaussian disturbances, we injected temporally correlated AR(1)-Laplace noise (Fig. 2). Despite heavy tails and temporal correlation, we observed an SNR gain of $6.9$ dB and an average noise reduction of $54.4\%$, with no visible phase lag. While performance is lower than under Gaussian noise, as to be expected, the degradation remains gradual. Crucially, the low-rank Hankel structure proved distribution-agnostic i.e., SVHT adaptively trims singular values inflated by heavy-tailed noise requiring explicit Gaussian assumptions on the noise statistics.
Comparison with Extended Kalman Filter (EKF)-based denoising: For reference, we evaluated an EKF as a denoising pipeline under the same Gaussian disturbance conditions defined in Sec. V-A. When the process and measurement covariance matrices are carefully tuned using approximate knowledge of the noise moments, the EKF achieves moderate reconstruction quality, yielding an SNR of approximately $0.6$ dB. However, even in this tuned setting the EKF exhibits a noticeable phase delay (Fig. 3, top), which increases the point-wise reconstruction error used in the SNR calculation and may be problematic for downstream control pipelines that require temporally aligned state estimates for stable feedback and prediction. In contrast, the proposed Page/Hankel-based method achieves an SNR of $19.2$ dB under the same disturbance conditions.
In practical robotic systems, the true measurement noise statistics are typically unknown and may vary over time. Since the EKF relies on explicit specification of these noise moments to compute the Kalman gain, its performance is sensitive to covariance mismatch. Under mild mismatch in the assumed noise statistics, reconstruction quality yields an SNR of approximately $10.0$ dB (Fig. 3, bottom). Our proposed Page/Hankel framework estimates the noise structure directly from the data through low-rank structure and adaptive SVHT thresholding, avoiding explicit noise variance specification while maintaining consistent reconstruction performance across noise regimes.
Figure 3: Comparison of Page/Hankel denoising and EKF filtering under temporally correlated Gaussian noise (AR(1)). Top: EKF with tuned covariance matrices (call-outs: SNR +0.6+0.6 dB). Bottom: EKF under covariance mismatch (cal-outs: +10.0+10.0 dB).
@@ V-B System Hardware and Experiments
This section experimentally validates the complete method in a setup (Fig. 4) motivated by a naval application. Currently, payloads are loaded into their silos by a ship-board crane while the ship docks in a port, where no ocean swells affect the safety of the loading process. It would be desirable to load these payloads while the ship is underway, potentially in heavy seas. In active sea states, the ship sways in response to ocean swells. Anticipating deck motion via a learned predictive model would enable the crane controller to compensate during payload delivery.
The testbed is a scaled crane mounted on an E2M eMove eM6-300-1500 five-DoF Stewart platform that can reproduce wave-induced ship deck motions. A VN-100 IMU measures orientation (quaternion) and angular velocity of the simulated moving base at 30 Hz. These measurements correspond to the platform (deck) motion only, which is modeled as a time-varying base motion input to the crane-payload system. Using these IMU data, the framework generates a $31$-step horizon open-loop forecast ($\approx$1.0 s) of the platform motion. Payload and crane dynamics are not directly predicted and are instead assumed to evolve downstream in response to the predicted base motion.
@@
Given a fixed error-tolerance $\varepsilon=0.04$ [^8] The tolerance $\varepsilon$ is chosen based on prior operational constraints of the experimental platform. In earlier implementations using a baseline constant-velocity predictor, prediction errors up to $0.08$ m were acceptable for reliable payload insertion. We therefore set $\varepsilon=0.04$ as a conservative threshold that remains consistent with the hardware constraints., we define a violation-duration metric on the prediction error $e_{t}$ as
@@
Together, these results demonstrate that our framework not only denoises real sensor streams but also yields stable and bounded predictions that are suitable for downstream planning and control applications. A real-time demonstration of the moving-base prediction and MPC integration is shown in the supplementary video [32].
V-C Parameter Sensitivity and Computational Analysis
While Sec.V-B validates hardware performance using a selected configuration of the hyper-parameters ($N=250$, $L=10$, $J=20$), we now examine the sensitivity of the prediction accuracy and computational cost to the embedding length $L$ and the number of Cadzow iterations $J$.
TABLE I: Effect of Cadzow iterations $J$ on runtime and prediction error ($N=250$, $L=10$). Cadzow time corresponds to the iterative Cadzow projection up to a maximum of $J$ iterations, with early termination when the relative tolerance $\texttt{tol}=10^{-6}$ is satisfied.
NN $L$ $J$ Cadzow Time (ms) Total Time (ms) RMSE
250 10 1 $0.6772\pm 0.1673$ $2.3220\pm 0.5380$ $0.012901$
250 10 5 $2.2844\pm 0.0812$ $4.0004\pm 0.1495$ $0.012601$
250 10 10 $4.1299\pm 0.0746$ $5.8169\pm 0.1208$ $0.011727$
250 10 20 $7.9089\pm 0.0642$ $9.5817\pm 0.0825$ $0.011505$
250 10 30 $11.7770\pm 0.5206$ $13.4480\pm 0.5879$ $0.011609$
Cadzow iterations $J$: Based on Table I, with $N=250$ and $L=10$ fixed, increasing the number of Cadzow iterations improves prediction accuracy up to $J=20$, reducing RMSE from $0.0129$ ($J=1$) to $0.0115$ ($J=20$). Beyond this point, additional iterations provide negligible accuracy gains while increasing computational cost approximately linearly. These results indicate that moderate Cadzow iterations provide a favorable trade-off between accuracy and runtime.
Embedding length $L$: As evidenced in Table II with $N=250$ and $J=20$ fixed, the prediction accuracy improves as the $L$ increases up to an intermediate value. RMSE decreases from $0.060$ at $L=4$ to $0.0108$ at $L=12$, with a slight degradation at $L=13$. This reflects a finite-data trade-off where larger embeddings increase model expressivity but reduce the number of snapshot pairs of $N/L$, which may degrade the matrix conditioning and stability of the locally linear dynamical model. Runtime grows moderately with $L$, with the Cadzow projection constituting the dominant over the total computational time. A comprehensive ablation study analyzing the sensitivity of RMSE, Cadzow runtime, and total computation time to the hyper-parameters is summarized in the Appendix (Table III).
TABLE II: Effect of embedding length $L$ on total computation time and prediction error ($N=250$, $J=20$). Cadzow time corresponds to the iterative Cadzow projection up to a maximum of $J$ iterations, with early termination when the relative tolerance $\texttt{tol}=10^{-6}$ is satisfied.
NN $L$ $J$ Cadzow Time (ms) Total Time (ms) RMSE
250 4 20 $7.8140\pm 3.2639$ $9.3070\pm 4.8364$ $0.060471$
250 6 20 $8.1512\pm 1.8150$ $9.9204\pm 2.1636$ $0.017239$
250 8 20 $7.8828\pm 0.4858$ $9.5510\pm 0.5851$ $0.013038$
250 10 20 $7.9628\pm 0.5314$ $9.5880\pm 0.6338$ $0.011505$
250 12 20 $8.1490\pm 0.5678$ $9.8960\pm 0.6983$ $0.010766$
@@ VI Conclusion And Future Work
We introduced a Page–Hankel SVHT framework for learning short-horizon predictive models of moving obstacles from noisy data streams. Simulations (Gaussian and heavy-tailed noise) showed good variance-stable, distribution-agnostic performance, and hardware tests with IMU data produce reliable forecasts that can support downstream control. Future work will explore online adaptation, tensorized delayed embeddings for multi-axis consistency, and integration within an MPC-based prediction framework.